answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the shape of this distribution?

A company is studying the number of daily debit card purchases. If there were 20 purchases and the probability of a debit card purchase is 0.5

What is the shape of this distribution

What is the probability that a 3-digit positive integer picked at random will have one or more 7 in its digits?

A three digit number can be chosen in 9*10*10 = 900ways

No. of ways to choose a three digit number without 7 is 8*9*9 = 648 ways

probability of picking a three digit number that includes atleast one digit that is 7 is 1- (648/900) = 252/900 = 7/25

Why is sample size important in determining probability?

The bigger the sample size the more accurate the results will be.

For example, if you roll a 6 sided die and track the results to get the probability of rolling a six. If you only roll 6 times, then you may not even get one 6 or you could get a few. A small sample size means you won't get very reliable results.

What is the probability that you will guess the same number that the computer chose in one try?

The probability depends on the range of numbers from which to choose.

If there is no limit to the range then probability is incalculable.

What is the product rule and the sum rule of probability?

Sum Rule:

P(A) = \sum_{B} P(A,B)

Product Rule:

P(A , B) = P(A) P(B|A) or P(A, B)=P(B) P(A|B)

[P(A|B) means probability of A given that B has occurred]

P(A, B) = P(A) P(B) , if A and B are independent events.

The probability of getting heads twice and tails once if a coin is tossed 3 times?

3/8, or 37.5%

There are eight possible sequences that can emerge from three coin tosses. Three have heads twice and tails once. Three have the opposite, two tails and one heads. The remaining possibilities are all heads or all tails.

  1. H H T
  2. H T H
  3. T H H
  4. T T H
  5. T H T
  6. H T T
  7. H H H
  8. T T T

Why can't the probability of an event occurring be greater than 1?

If probability becomes greater than one, then the uncertainty coefficient fluctuates and results in a terminal case.

If two fair dice are rolled what is the probability of rolling a double given that the sum was 11?

Let A = rolling a double Let B = sum is 11 P(A)=6/36=1/6 P(B)=2/36=1/18 since (5,6) and (6,5) produce a sum of 11. We want to find P(A/B)= P(A & B) / P(B) = 0 / P(B)=0 P(A & B) represent the event getting a double and the sum being 11.

What are the chances for dissipation of small amounts of fluid around lungs?

If there is a small pleural effusion the treatment of the underlying cause will reverse the process allowing the fluid to recede back into the parenchyma of the lungs. Normally this is the case with CHF patients that have pulmonary edema. Right sided heart failure causes a fluid buildup that can go as far as the pleural space.

From a pack of 52 cards red face cards are removed After that a card is drawn at random from the pack Find the probability that the card drawn is you a queen ii a red card iii a spade card?

Six face cards have been removed leaving 46 cards. 1) There are 2 quees left. P( queen) = 2/46 = 1/23 2)There are 20 red cards left. P( red card)=20/46=10/23 3) There are 26 spade cards left. P( spade card)=26/46 =13/23

How do you work out the mean in maths?

add up all the numbers you have then divide by the amount of numbers you have

The mean is the same as the average. You must add up all the data, then divide by the number of data entries you added. E.g. 3+5+2+6+4=20 then 20 divided by 5 = 4, so in this case 4 is the mean.

If a fair six sided die was rolled 4 times what is the probability of getting at least one ace one?

p(one) = 1/6

p(not one) = 5/6

p(at least one) = 1 - p(failure)

p(failure) = (4C0)(5/6)^4 = 1 x 625/1,296 = 0.48

p(at least one) = 1 - 0.48 = 0.52

How do you teach probability of independent events?

Two events are said to be independent if the result of the second event is not affected by the result of the first event. Some common ways to teach this are to perform simulations with coin flips.

Students need to understand that if A and B are independent events, the probability of both events occurring is the product of the probabilities of the individual events.

Students can predict and then observe probabilities of a fixed number of heads or tails.

This lets then see the ideas in action.

How do you use square roots on your TI-30X IIS calculator?

Well i had the same qustion first u press 2nd then you press the button that looks like this ^ then put your number inexample: 3 2nd ^ 78 it would be 4.272658682

A spinner has 13 equal sized spaces It has 4 red spaces 4 green spaces 3 yellow spaces and 2 orange spaces If Jake spins the spinner 2 times what is the probability he will get an orange space o?

all that matters in this problem, is that 2 of the 13 spaces are orange.

On the first spin, the odds are 11:2 against Jake getting the orange space.

On the second spin, the result of the first spin has no relevance or influence at all, so the odds are still 11:2 against.

Why is this question flagging up in "breakups"?

Trending Questions
. The bank earns 10 percent on commercial loans and 12 percent on home loans. What amount of money should be allotted for each type of loan to maximize the interest income? If two twelve side fair dice are rolled what it the probability that the total number of spots shown is equal to 5? What is strandard deviation? There are slips of paper that are numbered from 1 to 15 in a hat. What is the theoretical probability that a slip of paper that is chosen will be an even number? The likelihood that an event will occur like tossing a coin? A deck of cards is shuffled. What is the probability that all spades are together? What is the probability of pulling a red marbles out of a bag containing 16 red marbles 10 green marbles 14 blue marbles and 10 yellow marbles? What is the probability that all four 1 seeded teams make it to the Final Four? What is a probability of a not event? When two dice are rolled find the probability of not rolling a sum of 5? What is the probability that this couple will have a colorblind child? If a woman has given birth to two sons what is the probability that her third child will be a girl? What is the duce in a pack of playing cards? What is a difference between F statistics F distribution? What is probability of 1 even die and 1 odd when rolling 2 dice? You roll two dice What is the probability of the event that you roll a total of seven or twelve on the two dice? How can you be sure of getting the expected 5050 ratio from flipping a coin How are the face of a coin similar to genes? What is the probability if you roll a die and it lands on 5 and you flip a coin and it lands on heads? Differentiate between Worst case analysis and Monte Carlo analysis? How many different sums of money can be made using at most one bill of each of the folloing denominations 1 5 10 20 50?